OpenStudy (anonymous):

Brooke worked 10 hours more than Gregory last month. If Brooke worked 9 hours for every 2 hours that Gregory worked, how many hours did they each work?

ganeshie8 (ganeshie8):

Let : \(b\) = number of hours Broke worked \(g\) = number of hours Greg worked

ganeshie8 (ganeshie8):

Try setting up two equations based on the given info

OpenStudy (anonymous):

so what would the equation look like

ganeshie8 (ganeshie8):

`Brooke worked 10 hours more than Gregory last month.` translates to \[b = g+10\]

ganeshie8 (ganeshie8):

`Brooke worked 9 hours for every 2 hours that Gregory worked,` translates to \[\dfrac{b}{9}=\dfrac{g}{2}\]

ganeshie8 (ganeshie8):

two equations and two unknowns you should be able to solve